OpenStudy (anonymous):

What is the concentration of a solution of HNO3 if 200 mL of an 8M stock solution of the acid was diluted to 500 mL? How much water was added to the 200 mL volume when is was diluted?

OpenStudy (mos1635):

 n1 = n2 C1*V1 = C2*V2 .......C2=..... V=V2-V1